Given the following, find the number of moles:14.2g of Co2





A. 0.32





B. 0.32 mol CO2





C. 0.4 mol CO2





D. 1.00 mol CO2

1
Expert's answer
2022-04-11T13:38:04-0400

n = m / M

M (CO2) = 40 g/mol

n (CO2) = 14.2 / 40 = 0.355 mol = 0.4 mol


C. 0.4 mol CO2